Determine whether (2, -3) is a solution to the following system of equations. Be sure to explain your answer.
3x - 2y = 12
2x + 3y = -5.... I am not sure how to do this, I tried graphing it and seeing what i could do from there but im just not getting it!..
AND...
When solving a system of equations using the substitution menthod, you obtain the following result: 8 = 9. This result gives you some information about the solution.
a) How many solutions does the system have?
b) What would the graph of the system look like?
I just do not even know how to do this one, I dont know where to even start!HELP!! ...

There are the three possible outcomes for a system of equations. What does each LOOK like? How many solutions does each one have?

Found 2 solutions by jennywecas, MathTherapy:
Answer by jennywecas(28)

Determine whether (2, -3) is a solution to the following system of equations. Be sure to explain your answer.
3x - 2y = 12
2x + 3y = -5....
First what you do is substitute 2 with x and -3 with y. the first equation is 3(2)-(-3)=12. 3*2=6. 6-(-3)=9 9 does not equal 12 so you know it has no solution.

When solving a system of equations using the substitution menthod, you obtain the following result: 8 = 9. This result gives you some information about the solution.
a) How many solutions does the system have? No solution
b) What would the graph of the system look like? the lines do ot touch at all


on the question "There are the three possible outcomes for a system of equations. What does each LOOK like? How many solutions does each one have?" i can't help you i am sorry.

Answer by MathTherapy(10552)


Determine whether (2, -3) is a solution to the following system of equations. Be sure to explain your answer.

3x - 2y = 12
2x + 3y = -5

Substitute the coordinate point (2, - 3) into the equation to see if it makes it true, with 2 being the value of x, and - 3 being the value of y

3x - 2y = 12

3(2) - 2(- 3) = 12

6 + 6 = 12 (TRUE)


2x + 3y = - 5

2(2) + 3(- 3) = - 5

4 - 9 = - 5 (TRUE)

Therefore, (2, - 3) is a solution to the system